x=y/z+1.00
sove for z algebraically.

Answer by Alan3354(69443)

x=y/z+1.00
sove for z algebraically.
----------
x=y/z+1
Subtract 1
x-1 = y/z
Multiply by z
z*(x-1) = y
Divide
z = y/(x-1)